In how many ways can 6 boys and 5 girls be seated in a circular manner?
Answer by ikleyn(52946)

.

In (6+5-1)! = (11-1)! = 10! = 10*9*8*7*6*5*4*3*2*1 = 3628800  different ways.

Solved.


-----------------

To see other similar solved problems,  look into the lesson
    - Persons sitting around a cicular table
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Combinatorics: Combinations and permutations".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.